OpenStudy (anonymous):

A gas occupies 200 mL at -73.0°C. To have the same gas occupy 300 mL, should the temperature be increased or decreased? What is the new temperature in Kelvins? K

thomaster (thomaster):

\[\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}}\] Solve this with the ideal gas law: \(\Large PV=nRT\) P= pressure (not given in your question so we assume it's 1 atm) V= volume n= number moles R= gas constant T= temperature in kelvin \(\Large\dfrac{V}{T}=\dfrac{nR}{P}\) \(\Large\dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}\) \(\Large\dfrac{200}{200}=\dfrac{300}{??}\) \(\Large \dfrac{200*300}{200}=300K=27°C\)